OpenStudy (purplemexican):

Laura has moved to a new apartment. Her schoolbooks comprising of different subjects are mixed in a bag during the move. Four books are of mathematics, three are English, and six are science. If Laura opens the bag and selects books at random, find the given probability. P(3 English books)

Vocaloid (vocaloid):

total number of books = 4 + 3 + 6 = 13 probability of the first English book: (3/13) probability of the second English book: (2/12), since we already picked one probability of the third English book: (1/11) probability = (3/13)*(2/12)*(1/11)